# SPLITZ: Certifiable Robustness via Split Lipschitz Randomized Smoothing thanks: This work was supported by NSF grants CCF 2100013, CNS 2209951, CCF 1651492, CNS 2317192 and CNS 1822071. This wrok has been submitted to IEEE Transactions on Information Forensics and Security

### v1

@@ Abstract
Certifiable robustness gives the guarantee that small perturbations around an input to a classifier will not change the prediction. There are two approaches to provide certifiable robustness to adversarial examples– a) explicitly training classifiers with small Lipschitz constants, and b) Randomized smoothing, which adds random noise to the input to create a smooth classifier. We propose SPLITZ, a practical and novel approach which leverages the synergistic benefits of both the above ideas into a single framework. Our main idea is to split a classifier into two halves, constrain the Lipschitz constant of the first half, and smooth the second half via randomization. Motivation for SPLITZ comes from the observation that many standard deep networks exhibit heterogeneity in Lipschitz constants across layers. SPLITZ can exploit this heterogeneity while inheriting the scalability of randomized smoothing. We present a principled approach to train SPLITZ and provide theoretical analysis to derive certified robustness guarantees during inference. We present a comprehensive comparison of robustness-accuracy tradeoffs and show that SPLITZ consistently improves upon existing state-of-the-art approaches on MNIST and CIFAR-10 datasets. For instance, with $\ell_{2}$ norm perturbation budget of $\epsilon=1$, SPLITZ achieves 43.2% top-1 test accuracy on CIFAR-10 dataset compared to state-of-art top-1 test accuracy 39.8%.
@@ I Introduction
As deep learning becomes dominant in many important areas, ensuring robustness during test time becomes increasingly important. Deep neural networks are known to be vulnerable to adversarial attacks: small imperceptible perturbations in the inputs leading to incorrect decisions [1, 2]. Although many works have proposed heuristic defenses for training robust classifiers, they are often shown to be inadequate against adaptive attacks [3, 4, 5]. Therefore, a growing literature on certifiable robustness has emerged [6, 7, 8]; where the classifier’s prediction must be provably robust around any input within a perturbation budget.
@@
Ensuring certified robustness by constraining the Lipschitz constant usually involves the estimation of the Lipschitz constant for an arbitrary neural network. The main challenge is that accurate estimation of Lipschitz constants becomes hard for larger networks, and upper bounds become loose leading to vacuous bounds on certified radius. Recently, a variety of approaches have emerged that focus on training while explicitly constraining the model’s Lipschitz constant (outputs), for instance, LMT [18] and BCP [11]. These methods control the Lischitz constant of the model by constraining the outputs of each layer (or the outputs of the model); this has the dual benefit of enhanced robustness of the model as well as better estimation of the overall Lipschitz constant of the model. To further minimize the Lipschitz constant during the certified robust training process and better estimate the local Lipschitz constant of the model, recent work [19] focus on the constrained training with respect to the local Lipschitz constant by utilizing the clipped version of the activation functions.
@@
Lipschitz constrained training provides deterministic guarantees on certified radius and is often challenging to accurately estimate the Lipschitz constant of a large neural network. RS on the other hand offers scalability to arbitrarily large networks and provide the closed-form certified robust radius. These guarantees, however, are probabilistic in nature and the smoothing procedure treats the entire classifier as a black box mechanism.
TABLE I: Comparison of certified test accuracy (%) on MNIST under $\ell_{2}$ norm perturbation. Each entry lists the certified accuracy using numbers taken from respective papers (RS results follow from previous benchmark papers [22, 12]).
@@
Method Extra data 1.50 1.75 2.00 2.25 2.50
RS [7] 67.3 46.2 32.5 19.7 10.9
MACER [16] 73.0 50.0 36.0 28.0 -
Consistency [22] 82.2 70.5 45.5 37.2 28.0
SmoothMix [12] 81.8 70.7 44.9 37.1 29.3
DRT [23] 83.3 69.6 48.3 40.3 34.8
RS+OrthoNN [21] 70.1 49.7 33.2 21.0 10.5
SPLITZ 80.2 71.3 62.3 51.7 40.5
@@
Overview of SPLITZ and Contributions. In this paper, we propose SPLITZ, which combines and leverages the synergies offered by both local Lipschitz constrained training and randomized smoothing. The general idea is to split a classifier into two halves: the first half (usually a few layers) is constrained to keep a smaller Lipschitz (upper) bound, and the latter half of the network is smoothed via randomization. We propose the use of local Lipschitz constant(s) of the first half of the network. This is because it can capture the stability of the model with respect to each individual input. To enhance certified robustness, we incorporated the local Lipschitz constant of the network’s first half into the loss function as a regularization term. This approach aims to maintain a comparatively small local Lipschitz constant (typically less than 1) for the network’s first half, thereby improving the certified robustness. To obtain the certified radius during inference, we derive a closed-form expression for the certified radius based on the local Lipschitz constant as well as the randomized smoothing parameters as shown in Theorem 1; this result illustrates that enforcing a relatively small local Lipschitz constant can help in improving the certified radius. To the best of our knowledge, this is the first systematic framework to combine the ideas of local Lipschitz constants with randomized smoothing.
Interestingly, this approach yields state-of-the-art results for several datasets. For instance, Table I compares the certified test accuracy of SPLITZ and existing state-of-art techniques for various values of $\epsilon$ (perturbation budget or certified radius) on the MNIST dataset. For $\epsilon$ ($\ell_{2}$ norm perturbation size) even as large as $2.5$, where the state-of-the-art accuracy is $34.8\%$, SPLITZ achieves certified accuracy of around $40.5\%$. In Section IV, we present comprehensive set of results on MNIST and CIFAR-10 datasets. In addition, we also provide a detailed ablation study, and study the impact of various hyperparameters (such as the location of the split, randomized smoothing parameters).
@@
The paper is organized as follows: Section II introduces the objectives of the paper, preliminaries, and definitions of randomized smoothing and Lipschitz constant(s) used in the paper. Additionally, we review related works concerning randomized smoothing and Lipschitz training. Section III discusses the theoretical guarantees of SPLITZ and the corresponding training mechanisms. Experiments and evaluation results are presented and discussed in Section IV. Furthermore, additional experimental details on the Lipschitz constant, theoretical results on the local Lipschitz bound, and supplementary experimental results of the main findings are provided in the Appendices A, B, and C, respectively. Our code is available at: https://anonymous.4open.science/r/SPLITZ-3C66.
@@ II Preliminaries on Certified Robustness
$\mathbb{P}(f(x)=f(x^{\prime}))\geq 1-\alpha$, $\forall x^{\prime}$, such that $x^{\prime}=x+\delta,\parallel\delta\parallel_{p}\leq\epsilon$
@@ III SPLITZ: Inference, Certification and Training
In this Section, we first describe the details of the proposed SPLITZ classifier along with the motivation as well as key distinctions from prior work. We then present new theoretical results on certified radius for SPLITZ. Subsequently, we describe the training methodology for SPLITZ as well as inference and computation of the certified radius. Suppose we are given a base classifier $f:\mathbb{R}^{d}\rightarrow\mathcal{Y}$ which is a composition of $K$ functions. Consider an arbitrary “split” of $f$ as $f(\cdot)=f_{\text{R}}(f_{\text{L}}(\cdot))\triangleq f_{\text{R}}\circ f_{\text{L}}$. As an example, if the classifier has $K=2$ hidden layers, i.e., $f(x)=f_{2}(f_{1}(x))$, then there are $K+1=3$ possible compositions/splits:
$a)f_{\text{R}}=I,f_{\text{L}}=f_{2}\circ f_{1},$
$b)f_{\text{R}}=f_{2},f_{\text{L}}=f_{1},$
$c)f_{\text{R}}=f_{2}\circ f_{1},f_{\text{L}}=I,$
where $I$ represents the identity function.
@@
$g_{\textit{SPLITZ}}(x)=\underset{c~{}\in~{}\mathcal{Y}}{argmax}~{}\underset{\delta\sim\mathcal{N}(0,\sigma^{2}I)}{\mathbb{P}}(f_{\text{R}}(f_{\text{L}}(x)+\delta)=c)$ (5)
@@
Figure 3: Comparison of certified radius with $\ell_{2}$ norm perturbation w.r.t RS [7] and SPLITZ (ours), when varying levels of Gaussian noise $\sigma\in\{0.25,0.5,1.0\}$ on the CIFAR-10 dataset. We can observe that SPLITZ consistently outperforms RS [7] under different noise levels. For instance, when Radius = 1.0, SPLITZ achieves $43.2\%$ certified test accuracy while the best certified test accuracy of RS under different noise levels is $22.0\%$. We refer the reader to Table II for a comprehensive comparison of SPLITZ with several other approaches on CIFAR-10 dataset.
Given an input $x$, in order to compute the certified radius for SPLITZ classifier, we need $L_{f_{\text{L}}}^{(\gamma)}(x)$, i.e., the $\gamma$-local Lipschitz constant (discussed in the next Section) and the certified radius of right half of the classifier, i.e., $R_{f_{\text{R}}}(f_{\text{L}}(x))$. For Gaussian noise perturbation in the second half, $R_{f_{\text{R}}}(f_{\text{L}}(x))$ is exactly the randomized smoothing $\ell_{2}$ radius [7], given as $R_{f_{\text{R}}}(f_{\text{L}}(x))=\frac{\sigma}{2}(\Phi^{-1}(\underline{p_{A}})-\Phi^{-1}(\overline{p_{B}}))$, where $\underline{p_{A}}$ is the lower bound of the probability that the most probable class $c_{A}$ is returned, $\underline{p_{B}}$ is upper bound of the probability that the “runner-up” class $c_{B}$ is returned.
@@
Optimization over $\gamma$ We note from Theorem 1 that finding the optimal choice of $\gamma$ is crucial. One way is to apply the efficient binary search during the certify process to find the optimal value of $\gamma$. Specifically, we set the initial value of $\gamma$ and compute the corresponding local Lipschitz constant $L_{f_{L}}^{(\gamma)}(x)$ at input $x$. By comparing the value between $\gamma$ and ${R_{f_{R}}(f_{L}(x))}/{L_{f_{L}}^{(\gamma)}(x)}$, we divide the search space into two halves at each iteration to narrow down the search space until $\gamma^{*}={R_{f_{R}}(f_{L}(x))}/{L_{f_{L}}^{(\gamma^{*})}(x)}$. Another way is to do a one-step search. Specifically, we first approximate the local Lipschitz constant $\tilde{L}_{f_{L}}^{(\gamma)}(x)$ at $x$ by averaging local Lipschitz constants of inference data given the inference $\gamma$. We then set $\gamma^{\prime}={R_{f_{R}}(f_{L}(x))}/{\tilde{L}_{f_{L}}^{(\gamma)}(x)}$ and re-calculate the local Lipschitz constant $\tilde{L}_{f_{L}}^{(\gamma^{\prime})}(x)$ according to $\gamma^{\prime}$. Finally, we compute the approximate optimal $\gamma^{*}={R_{f_{R}}(f_{L}(x))}/{\tilde{L}_{f_{L}}^{(\gamma^{\prime})}(x)}$. Overall, we show the certification process in detail in Algorithm 1.
@@
Compatibility of SPLITZ with other defenses In addition, we claim that the SPLITZ mechanism is also compatible with other RS based certified robust techniques, such as adversarial smoothing [15], mixsmoothing [12] or denoising diffusion models [14]. As an example, [14] propose a denosing mechanism using a diffusion model, which achieves the state-of-the-art. Our SPLITZ classifier contains two parts, left part is constrained by a small local Lipschitz constant while right part is smoothed by noise, which is same as a randomized smoothing based mechanism. Thus, our model can easily add a diffusion denoising model after the noise layer (after $f_{L}(x)+\delta$) and then feed the denoised samples into the $f_{R}$. Similarly, for adversarial smoothing or mixsmoothing, SPLITZ is adaptable to feed either adversarial examples ($f_{L}(x^{\prime})+\delta$) or mixup samples ($f_{L}(\tilde{x})+\delta$) respectively to the right half of the classifier $f_{R}$. On the other hand, SPLITZ is also compatible to other Lipschitz constrained based mechanisms (e.g., orthogonal neural network based mechanism [21, 20]). For instance, we can incorporate an orthogonal constraint regularization into the loss function to enforce orthogonality within the convolution layer throughout the training process. The outcomes of integrating the orthogonal convolution neural network with SPLITZ are meticulously detailed in the Section IV.
@@
We first note from Definition 3 that $g_{\textit{SPLITZ}}(x)$ can be written as $g_{\textit{SPLITZ}}(x)=\tilde{g}(f_{L}(x))$, where the function $\tilde{g}$ is the smoothed version of $f_{R}$.
We are given that the smooth version $\tilde{g}$ has a certified radius of $R_{\tilde{g}}(u)\triangleq R_{f_{\text{R}}}(u)$ with probability at least $1-\alpha$. This is equivalent to the statement that for all $u^{\prime}$ such that $||u-u^{\prime}||_{p}\leq R_{f_{\text{R}}}(u)$, we have:
@@
$L_{f_{L}}^{(\gamma)}(x)||x-x^{\prime}||_{p}\leq R_{f_{\text{R}}}(f_{\text{L}}(x))$
$\leftrightarrow||x-x^{\prime}||_{p}\leq\frac{R_{f_{R}}(f_{L}(x))}{L_{f_{L}}^{(\gamma)}(x)}$ (11)
@@
where $\lambda\in[0,1]$ is a hyperparameter controlling the tradeoff between the accuracy and robustness, $\theta$ is a learnable parameter to optimize the local Lipschitz constant, and $\text{Loss}(\cdot)$ is the loss function (e.g., cross entropy loss). Following the literature on randomized smoothing, we replace the expectation operators with their empirical estimates, and our loss function becomes:
@@
where $W_{s}$ is the weight matrix of layer $s$.
@@
  Input: Training set $D_{train}=\{x_{i},y_{i}\}_{i=1}^{N_{train}}$; noise level $\sigma$, training steps $T$, Lipschitz threshold $\theta$, training hyperparameter $\gamma$.
@@
Summary of SPLITZ Training Methodology Overall, our training procedure is presented in Algorithm 2. During the process of computing local Lipschitz constant of $f_{L}$, for each iteration, we feed the input to the classifier $f_{L}$ and calculate the LB and UB of outputs of each affine layer in $f_{L}$ given the input $x$ within a $\gamma$ ball. We then can calculate the indicator matrix $I^{v}$ and compute the spectral norm of the reduced weight matrix $\parallel I^{v}_{s}W_{s}I^{v}_{s}\parallel$ for each layer $s$ in $f_{L}$ using power iteration. By multiplying the reduced weight matrix of each affine layer in $f_{L}$, we are able to arrive at the local Lipschitz constant of $f_{L}$. Secondly, we smooth the right half of the neural network $f_{R}$ by sampling from Gaussian noise with zero mean and adding it to the output of $f_{L}$. Then we feed the noisy samples $f_{L}(x)+\delta$ to $f_{R}$ and obtain the corresponding loss. Next, we minimize the overall loss and backward the parameters to optimize the overall network $f$. Finally, we certify the base classifier $f$ to obtain the Lipschitz smoothing classifier $g_{\textit{SPLITZ}}$ as shown in Algorithm 1.

### v2

# SPLITZ: Certifiable Robustness via Split Lipschitz Randomized Smoothing thanks: This work was supported by NSF grants CCF 2100013, CNS 2209951, CCF 1651492, CNS 2317192 and CNS 1822071. This work has been submitted to IEEE Transactions on Information Forensics and Security

@@ Abstract
Certifiable robustness gives the guarantee that small perturbations around an input to a classifier will not change the prediction. There are two approaches to provide certifiable robustness to adversarial examples– a) explicitly training classifiers with small Lipschitz constants, and b) Randomized smoothing, which adds random noise to the input to create a smooth classifier. We propose SPLITZ, a practical and novel approach which leverages the synergistic benefits of both the above ideas into a single framework. Our main idea is to split a classifier into two halves, constrain the Lipschitz constant of the first half, and smooth the second half via randomization. Motivation for SPLITZ comes from the observation that many standard deep networks exhibit heterogeneity in Lipschitz constants across layers. SPLITZ can exploit this heterogeneity while inheriting the scalability of randomized smoothing. We present a principled approach to train SPLITZ and provide theoretical analysis to derive certified robustness guarantees during inference. We present a comprehensive comparison of robustness-accuracy trade-offs and show that SPLITZ consistently improves on existing state-of-the-art approaches in the MNIST, CIFAR-10 and ImageNet datasets. For instance, with $\ell_{2}$ norm perturbation budget of $\epsilon=1$, SPLITZ achieves 43.2% top-1 test accuracy on CIFAR-10 dataset compared to state-of-art top-1 test accuracy 39.8%.
@@ I Introduction
As deep learning becomes dominant in many important areas, ensuring robustness becomes increasingly important. Deep neural networks are known to be vulnerable to adversarial attacks: small imperceptible perturbations in the inputs leading to incorrect decisions [1, 2]. Although many works have proposed heuristic defenses for training robust classifiers, they are often shown to be inadequate against adaptive attacks [3, 4, 5]. Therefore, a growing literature on certifiable robustness has emerged [6, 7, 8]; where the classifier’s prediction must be provably robust around any input within a perturbation budget.
@@
Ensuring certified robustness by constraining the Lipschitz constant usually involves estimating the Lipschitz constant for an arbitrary neural network. The main challenge is that accurate estimation of Lipschitz constants becomes hard for larger networks, and upper bounds become loose leading to vacuous bounds on certified radius. Thus, a variety of approaches have emerged that focus on training while explicitly constraining the model’s Lipschitz constant (outputs), for instance, LMT [18] and BCP [11]. These methods control the Lischitz constant of the model by constraining the outputs of each layer (or the outputs of the model); this has the dual benefit of enhanced robustness of the model as well as better estimation of the overall Lipschitz constant of the model. To further minimize the Lipschitz constant during the certified robust training process and better estimate the local Lipschitz constant of the model, recent work [19] focus on the constrained training with respect to the local Lipschitz constant by utilizing the clipped version of the activation functions.
@@
Lipschitz constrained training provides deterministic guarantees on certified radius and is often challenging to accurately estimate the Lipschitz constant of a large neural network. RS on the other hand offers scalability to arbitrarily large networks and provide the closed-form certified robust radius. These guarantees, however, are probabilistic in nature and the smoothing procedure treats the entire classifier as a black box.
@@
MNIST CIFAR-10 ImageNet
Method 1.50 1.75 2.00 0.5 0.75 1.0 1.5 2.0 3.0
RS [7] 67.3 46.2 32.5 43.0 32.0 22.0 29.0 19.0 12.0
MACER [16] 73.0 50.0 36.0 59.0 46.0 38.0 31.0 25.0 14.0
Consistency [22] 82.2 70.5 45.5 58.1 48.5 37.8 34.0 24.0 17.0
SmoothMix [12] 81.8 70.7 44.9 57.9 47.7 37.2 38.0 26.0 20.0
DRT [23] 83.3 69.6 48.3 60.2 50.5 39.8 39.8 30.4 23.2
RS+OrthoNN [21] 70.1 49.7 33.2 45.9 28.9 19.2 28.4 16.2 10.0
SPLITZ 80.2 71.3 62.3 63.2 53.4 43.2 38.6 31.2 23.9
TABLE I: Comparison of certified test accuracy (%) on MNIST, CIFAR-10, and ImageNet under $\ell_{2}$ norm perturbation (see full results in Section IV). Each entry lists the certified accuracy using numbers taken from respective papers (RS results on MNIST follow from previous benchmark papers [22, 12]).
@@
Overview of SPLITZ and Contributions. In this paper, we propose SPLITZ, which combines and leverages the synergies offered by both local Lipschitz constrained training and randomized smoothing. The general idea is to split a classifier into two halves: the first half (usually a few layers) is constrained to keep a smaller Lipschitz (upper) bound, and the latter half of the network is smoothed via randomization. We propose the use of local Lipschitz constant(s) of the first half of the network. This is because it can capture the stability of the model with respect to each individual input. To enhance certified robustness, we incorporated the local Lipschitz constant of the network’s first half into the loss function as a regularization term. This approach aims to maintain a comparatively small local Lipschitz constant (typically less than 1) for the network’s first half, thereby improving the certified robustness. We provide the theoretical guarantee of SPLITZ and derive a closed-form certified radius based on the local Lipschitz constant as well as the randomized smoothing parameters as shown in Theorem 1; this result illustrates that enforcing a relatively small local Lipschitz constant can help in improving the certified radius. To the best of our knowledge, this is the first systematic framework to combine the ideas of local Lipschitz constants with randomized smoothing.
Interestingly, this approach yields state-of-the-art results for several datasets. For instance, Table I compares the certified test accuracy of SPLITZ and existing state-of-art techniques for various values of $\epsilon$ (perturbation budget or certified radius) on the MNIST, CIFAR-10 and ImageNet datasets. For $\epsilon$ as large as $2.0$, where the state-of-the-art accuracy is $48.3\%$, SPLITZ achieves certified accuracy of around $62.3\%$. In Section IV, we present comprehensive set of results on MNIST, CIFAR-10 and ImageNet datasets. In addition, we also provide a detailed ablation study, and study the impact of various hyperparameters (such as the location of the split, randomized smoothing parameters, effects of local Lipschitz constant).
@@
The paper is organized as follows: Section II introduces the objectives of the paper, preliminaries, and definitions of randomized smoothing and Lipschitz constant(s) used in the paper. Additionally, we review related works concerning randomized smoothing and Lipschitz training. Section III discusses the theoretical guarantees of SPLITZ and the corresponding training mechanisms. Experiments and evaluation results are presented and discussed in Section IV. Furthermore, additional experimental details on the Lipschitz constant, theoretical results on the local Lipschitz bound, and supplementary experimental results of the main findings are provided in the Appendices A, B, and C, respectively.
@@ II Preliminaries on Certified Robustness
$\mathbb{P}(f(x)=f(x^{\prime}))\geq 1-\alpha$, $\forall x^{\prime}$, such that $x^{\prime}=x+\delta,||\delta||_{p}\leq\epsilon$
@@ III SPLITZ: Inference, Certification and Training
In this Section, we first describe the details of the proposed SPLITZ classifier along with the motivation as well as key distinctions from prior work. We then present new theoretical results on certified radius for SPLITZ. Subsequently, we describe the training methodology for SPLITZ as well as inference and computation of the certified radius. Suppose we are given a base classifier $f:\mathbb{R}^{d}\rightarrow\mathcal{Y}$ which is a composition of $K$ functions. Consider an arbitrary “split” of $f$ as $f(\cdot)=f_{\text{R}}(f_{\text{L}}(\cdot))\triangleq f_{\text{R}}\circ f_{\text{L}}$. As an example, if the classifier has $K=2$ hidden layers, i.e., $f(x)=f_{2}(f_{1}(x))$, then there are $K+1=3$ possible compositions/splits: $a)f_{\text{R}}=I,f_{\text{L}}=f_{2}\circ f_{1},b)f_{\text{R}}=f_{2},f_{\text{L}}=f_{1},c)f_{\text{R}}=f_{2}\circ f_{1},f_{\text{L}}=I$, where $I$ represents the identity function.
@@
$g_{\textit{SPLITZ}}(x)=\underset{c~{}\in~{}\mathcal{Y}}{\arg\max}~{}\underset{\delta\sim\mathcal{N}(0,\sigma^{2}I)}{\mathbb{P}}(f_{\text{R}}(f_{\text{L}}(x)+\delta)=c)$ (5)
@@
Figure 3: Comparison of certified radius with $\ell_{2}$ norm perturbation w.r.t RS [7], SmoothAdv [15], Smoothmix [12] and SPLITZ (ours), when varying levels of Gaussian noise $\sigma\in\{0.25,0.5,1.0\}$ on the CIFAR-10 dataset. We can observe that SPLITZ consistently outperforms RS [7] under different noise levels. For instance, when Radius = 1.0, SPLITZ achieves $43.2\%$ certified test accuracy while the best certified test accuracy of RS under different noise levels is $22.0\%$. We refer the reader to Table III for a comprehensive comparison of SPLITZ with several other approaches on CIFAR-10 dataset.
Given an input $x$, to compute the certified radius for SPLITZ classifier, we need $L_{f_{\text{L}}}^{(\gamma)}(x)$, i.e., the $\gamma$-local Lipschitz constant (discussed in the next Section) and the certified radius of right half of the classifier, i.e., $R_{f_{\text{R}}}(f_{\text{L}}(x))$. For Gaussian noise perturbation in the second half, $R_{f_{\text{R}}}(f_{\text{L}}(x))$ is exactly the randomized smoothing $\ell_{2}$ radius [7], given as $R_{f_{\text{R}}}(f_{\text{L}}(x))=\frac{\sigma}{2}(\Phi^{-1}(\underline{p_{A}})-\Phi^{-1}(\overline{p_{B}}))$, where $\underline{p_{A}}$ is the lower bound of the probability that the most probable class $c_{A}$ is returned, $\underline{p_{B}}$ is upper bound of the probability that the “runner-up” class $c_{B}$ is returned.
@@
Optimization over $\gamma$ We note from Theorem 1 that finding the optimal choice of $\gamma$ is crucial. One way is to apply the efficient binary search during the certify process to find the optimal value of $\gamma$. Specifically, we set the initial value of $\gamma$ and compute the corresponding local Lipschitz constant $L_{f_{L}}^{(\gamma)}(x)$ at input $x$. By comparing the value between $\gamma$ and ${R_{f_{R}}(f_{L}(x))}/{L_{f_{L}}^{(\gamma)}(x)}$, we divide the search space into two halves at each iteration to narrow down the search space until $\gamma^{*}={R_{f_{R}}(f_{L}(x))}/{L_{f_{L}}^{(\gamma^{*})}(x)}$. Another way is to do a one-step search. Specifically, we first approximate the local Lipschitz constant $\tilde{L}_{f_{L}}^{(\gamma)}(x)$ at $x$ by averaging local Lipschitz constants of inference data given the inference $\gamma$. We then set $\gamma^{\prime}={R_{f_{R}}(f_{L}(x))}/{\tilde{L}_{f_{L}}^{(\gamma)}(x)}$ and re-calculate the local Lipschitz constant $\tilde{L}_{f_{L}}^{(\gamma^{\prime})}(x)$ according to $\gamma^{\prime}$. Finally, we compute the approximate optimal $\gamma^{*}={R_{f_{R}}(f_{L}(x))}/{\tilde{L}_{f_{L}}^{(\gamma^{\prime})}(x)}$. We study the behavior of the certified radius with respect to the value of $\gamma$ in Sec IV and we show that one step optimization is sufficient to achieve the desired certified radius with SPLITZ. Overall, we show the certification process in detail in Algorithm 1.
@@
Compatibility of SPLITZ with other defenses In addition, the SPLITZ mechanism is also compatible with other RS based certified robust techniques, such as adversarial smoothing [15], mixsmoothing [12] or denoising diffusion models [14]. For example, [14] propose a denosing mechanism using a diffusion model, which achieves the state-of-the-art. Our SPLITZ classifier contains two parts, left part is constrained by a small local Lipschitz constant while right part is smoothed by noise, which is same as a randomized smoothing based mechanism. Thus, our model can easily add a diffusion denoising model after the noise layer (after $f_{L}(x)+\delta$) and then feed the denoised samples into the $f_{R}$. Similarly, for adversarial smoothing or mixsmoothing, SPLITZ is adaptable to feed either adversarial examples ($f_{L}(x^{\prime})+\delta$) or mixup samples ($f_{L}(\tilde{x})+\delta$) respectively to the right half of the classifier $f_{R}$. On the other hand, SPLITZ is also compatible to other Lipschitz constrained based mechanisms (e.g., orthogonal neural network based mechanism [21, 20]). For instance, we can incorporate an orthogonal constraint regularization into the loss function to enforce orthogonality within the convolution layer throughout the training process. The results of integrating the orthogonal convolution neural network with SPLITZ are meticulously detailed in Section IV.
@@
We first note from Definition 3 that $g_{\textit{SPLITZ}}(x)$ can be written as $g_{\textit{SPLITZ}}(x)=\tilde{g}(f_{L}(x))$, where the function $\tilde{g}$ is the smoothed version of $f_{R}$. We are given that the smooth version $\tilde{g}$ has a certified radius of $R_{\tilde{g}}(u)\triangleq R_{f_{\text{R}}}(u)$ with probability at least $1-\alpha$. This is equivalent to the statement that for all $u^{\prime}$ such that $||u-u^{\prime}||_{p}\leq R_{f_{\text{R}}}(u)$, we have:
@@
$L_{f_{L}}^{(\gamma)}(x)||x-x^{\prime}||_{p}\leq R_{f_{\text{R}}}(f_{\text{L}}(x))\leftrightarrow||x-x^{\prime}||_{p}\leq\frac{R_{f_{R}}(f_{L}(x))}{L_{f_{L}}^{(\gamma)}(x)}$ (11)
@@
where $\lambda\in[0,1]$ is a hyperparameter controlling the tradeoff between accuracy and robustness, $\theta$ is a learnable parameter to optimize the local Lipschitz constant, and $\text{Loss}(\cdot)$ is the loss function (e.g., cross entropy loss). Following the literature on randomized smoothing, we replace the expectation operators with their empirical estimates, and our loss function is:
@@
where $W_{s}$ is the weight matrix of layer $s$. We include more details with respect to local Lipschitz bound in Appendix B.
@@
  Input: Training set $D_{\text{train}}=\{x_{i},y_{i}\}_{i=1}^{N_{\text{train}}}$; noise level $\sigma$, training steps $T$, Lipschitz threshold $\theta$, training hyperparameter $\gamma$.
@@
Summary of SPLITZ Training Methodology Overall, our training procedure is presented in Algorithm 2. During the process of computing local Lipschitz constant of $f_{L}$, for each iteration, we feed the input to the classifier $f_{L}$ and calculate the LB and UB of outputs of each affine layer in $f_{L}$ given the input $x$ within a $\gamma$ ball. We then can calculate the indicator matrix $I^{v}$ and compute the spectral norm of the reduced weight matrix $\parallel I^{v}_{s}W_{s}I^{v}_{s}\parallel$ for each layer $s$ in $f_{L}$ using power iteration. By multiplying the reduced weight matrix of each affine layer in $f_{L}$, we are able to arrive at the local Lipschitz constant of $f_{L}$. Secondly, we smooth the right half of the neural network $f_{R}$ by sampling from Gaussian noise with zero mean and adding it to the output of $f_{L}$. Then we feed the noisy samples $f_{L}(x)+\delta$ to $f_{R}$ and obtain the corresponding loss. Next, we minimize the overall loss and backward the parameters to optimize the overall network $f$. Finally, we certify the base classifier $f$ to obtain the classifier $g_{\textit{SPLITZ}}$ as shown in Algorithm 1.